width of a rectangle flower bed is 7 ft less than the length. the area is 18 ft squared. find the length and width
I have to show my work?

Answer by checkley77(12844)

W=L-7
A=LW
18=L(L-7)
18=L^2-7L
L^2-7L-18=0
(L-9)(L+2)
L-9=0
L=9 ANS. FOR THE LENGTH.
W=9-7=2 ANS. FOR THE WIDTH.
PROOF:
18=9*2
18=18